# Byzantine Attacks Exploiting Penalties in Ethereum PoS

### v1

@@ 4 Inactivity Leak
The Ethereum PoS blockchain strives for the continuous growth of the finalized chain. In consequence, the protocol incentivize validators to finalize blocks actively. In the absence of finalization, validators incur penalties.
@@ 4.3 Stake’s functions during an inactivity leak
Validators with these three behaviors experience a different evolution in their inactivity score: (a) Active validators have a constant inactivity score $I(t)=0$; (b) Semi-active validators’ inactivity score increases by 3 every two epochs ($+4-1$), their inactivity score on average at $t$ is $I(t)=3t/2$; (c) Inactive validators’ inactivity score increases by 4 every epochs $I(t)=4t$. Thanks to Equation 3 we can determine the evolution of the stake of each type of validators during an inactivity leak:
@@ 5.1 GST upper bound for Safety
Finality on both chains is achieved precisely at $4686$ epochs after the beginning of the inactivity leak. Adding an epoch is necessary after gaining 2/3 of active stake to finalize the preceding justified checkpoint. This finalization ends the inactivity leak which has lasted approximately 3 weeks. Any network partition lasting longer than 4686 epochs will result in a loss of Safety because of conflicting finalization. This is an upper bound for Safety on the duration on inactivity leak with only honest validators.
@@ 5.2 Upper bound decrease due to Byzantine validators
We study the inactivity leak with these conditions: (i) at the beginning, less than one third of the stake is held by Byzantine validators ($\beta_{0}=n_{\rm B}/n<1/3$), the rest is held by honest validators ($1-\beta_{0}=n_{\rm H}/n$); (ii) The network is asynchronous (before GST); and (iii) Byzantine validators control the network delay, allowing them to be active on both branches while preventing honest validators from even observing the branch on which they are not active. They can manipulate message delays of honest nodes to simulate a partition between honest validators.

### v2

@@ 4 Inactivity Leak
The Ethereum PoS blockchain strives for the continuous growth of the finalized chain. In consequence, the protocol incentivizes validators to finalize blocks actively. In the absence of finalization, validators incur penalties.
@@ 4.3 Stake’s functions during an inactivity leak
Validators with these three behaviors experience a different evolution in their inactivity score: (a) Active validators have a constant inactivity score $I(t)=0$; (b) Semi-active validators’ inactivity score increases by 3 every two epochs ($+4-1$), their inactivity score on average at $t$ is $I(t)=3t/2$; (c) Inactive validators’ inactivity score increases by 4 every epoch $I(t)=4t$. Thanks to Equation 3, we can determine the evolution of the stake of each type of validator during an inactivity leak:
@@ 5.1 GST upper bound for Safety
Finality on both chains is achieved precisely at $4686$ epochs after the beginning of the inactivity leak. Adding an epoch is necessary after gaining 2/3 of active stake to finalize the preceding justified checkpoint. This finalization ends the inactivity leak which has lasted approximately 3 weeks. Any network partition lasting longer than 4686 epochs will result in a loss of Safety because of conflicting finalization. This is an upper bound for Safety on the duration of the inactivity leak with only honest validators.
@@ 5.2 Upper bound decrease due to Byzantine validators
We study the inactivity leak with these conditions: (i) at the beginning, less than one third of the stake is held by Byzantine validators ($\beta_{0}=n_{\rm B}/n<1/3$), the rest is held by honest validators ($1-\beta_{0}=n_{\rm H}/n$); (ii) The network is asynchronous (before GST); and (iii) Byzantine validators are not affected by network partitions.111111 In a model without partitions, one needs to give Byzantine validators more power to recreate our scenario. They must be able to control the network delay to allow them to be active on both branches while preventing honest validators from even observing the branch on which they are not active. They can manipulate message delays between groups of honest validators to simulate a partition between them.